OpenStudy (anonymous):

What are the directrices of the ellipse given by the equation x squared over 36 plus y squared over 100 = 1? @IMStuck

OpenStudy (jdoe0001):

hint: \(\bf \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\qquad directrix\to \cfrac{a^2}{c}\)

OpenStudy (jdoe0001):

recall that for an ellipse, \(\bf c=\sqrt{a^2-b^2}\)

OpenStudy (anteater):

Does the fact that the major axis is vertical make a difference as to which are a and b? Would the directrices be horizontal rather than vertical?

OpenStudy (jdoe0001):

ahemm yes, "a" is always the bigger denominator... so for this case then \(\bf \cfrac{(x-h)^2}{b^2}+\cfrac{(y-k)^2}{a^2}=1\)

OpenStudy (anteater):

Ok, thanks! :)